FEL Upgrade Project Weekly Brief – October 7-11, 2002




Installation and commissioning activities are proceeding. Good progress was made this week toward bringing the magnet trim power supplies and their controls up for the injector magnets. The EPICS control page for the trim supplies and the vacuum control page are up and functional in the Control Room. The injector magnets have been wired up. The Gun High Voltage Power Supply was tested to full specification (10.1 mA). This week the optical cavity (OC) vacuum chambers for the IR Upgrade were shipped from the vendor with delivery to JLab expected early next week. The OC chambers for the UV FEL Upgrade will follow by the end of the month.


On Friday Oct. 11, George Neil and Fred Dylla attended a briefing to the Secretary of Navy, Gordon England, presented by NAVSEA and ONR concerning the reconstitution of the Navy's High Energy Laser program that would include continued FEL development based on JLab technology. The Secretary characterized the proposed FEL development effort as "important and we [the Navy] should do it".

We held a Commissioning Planning Meeting this week to establish operations ground rules as we bring up the hardware. (See details below under the Operations Section).

WBS 4 (Injector):

The NEG pumps were installed in the gun chamber. When the gun chamber was pumped down a leak was discovered in a weld of a flange, which could easily be locally rewelded. The tapered shielded bellows and BPM/shielded bellows manufacturing was completed, these mount on either side of the lightbox. The stalk bellows/drive has had its parts cleaned and is awaiting welding. We decided to polish the ball cathode to only 9um prior to implanting. The decision was based on data from the implanted test electrodes which had been polished to that level and stood off 30MV/m. The ball cathode is currently at 600 grit finish with most of the etch pit damaged removed, 9um is the next step.

Gun HVPS - The HVPS electronics was modified to reach higher current. The unit was tested and achieved 10.1 ma. The FEL system specification calls for 10 ma from the Gun. Parts to combine the Running and Conditioning Resistors are being assembled. Several parts have still not been received, but are due within 2 weeks.



WBS 6 (RF):

Quarter HVPS - Work continues on the unit 4 klystron cart. It is needed before the interlocks can be checked and the 480 VAC wired to the unit. Differential input cards for the PLC were installed and tested this week on unit 3. This eliminated several isolation cards that were loading down numerous analog signals. Unit 4 is also being updated with these new cards.

Zone 4 - The special length waveguides for this zone are being fabricated and are due back late next week. It will take 1 or 2 days to install and test for leaks.

A TOSP is being written to allow the start of SRF commissioning without having the concrete blocks back in place at the roll-up door for a radiation shield. Portable lead sheets, a radiation monitor connected to the PSS, and low power pulsed beam will be used to mitigate the hazard.

WBS 8 (Instrumentation):

Software update: The beam viewer controls have been complete (for months), the magnet screens are being worked, the injector is complete as well as the injector power supply (these are running on upgraded serial communications to eliminate CAMAC), RF software is ready for commissioning, vacuum screens and controls are complete. Work continues on the Linux IOC for video acquisition, and the DSP based on line FEL spectrometer.

BLMs (Beam Loss Monitors) - The prototype cards are done. Their integration into EPICS is the primary focus of the project. The 3 remaining front cards (F0151) are with the EES group for assembly. Another MPS 12 Channel Beam Loss Monitor Breakout Module has been populated and is ready for installation. - Drive Laser Controls - restoring the variable attenuator and MPS shutter controls has shown to more involved than was originally expected. The variable attenuator has required that a full 4-channel motor control system be implemented in the drive laser clean room (to keep all drive laser controls together). The motor control system is currently in place and is being checked-out in EPICS. This system will also serve the future second drive laser that has been proposed. The variable attenuator can be seen at

http://laser.jlab.org/systems/dlpc/dlpc_var-atten_FL06_2002-10-11.jpg. The MPS shutter is centrally controlled by an interface chassis in the FEL control room (FL07B11) which needs to be modified to account for the new control method using the AcroMag card. The modifications to this chassis are underway.

Hardware upgrades of the trim magnets racks in FL05B09and B10 have been completed. Utility chassis modifications for magnet racks FL05B10 through13 have been completed and installed. Installation of the trim magnets on the beam line has begun in the 0F region. Progress was made on the Injection Dipole 55 kW PS. Testing of communications, hardware and software for the magnet trim rack in the 0F region (FL05B13) proved successful. Trim magnets MGJ0F04 and MQJ0F05 were powered and current ramped on these magnets.

Beam viewer Camera Assembly parts have been inventoried and are being assembled. The 5 wiggler viewers are ready for installation early next week. The installation will resume next week for the 1F and 2F region. Beam viewer signal & power cables have been terminated for the 0F, 1F and 2F regions (3F complete). Assembly of the viewers continues in these 2 areas. More Beam viewer Interface boxes are being assembled to support this effort.

The new gun HV monitor and interlock chassis is complete and testing will begin on Monday. This includes the connections to the Acromag VME module. Progress continues on the software for the Acromag module, these VME boards are used both for drive laser control and for the gun. We met with EECAD to determine the status of several "inactive" drawings in their queue. The drawings in question are still required to be completed, but are low on the priority list. The first of the new MPS boards is complete and programming/testing has begun. The balance of 4 are in fabrication. Four (4) 64 bit GP Interface modules were populated for the MPS system upgrade and are ready for testing. The 15 pin fanout boards are in house from the vendor. The injector BPM has been welded and tested good. This goes between the light box and buncher. Work continues to bring up the full video distribution system.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• MEI received most of the re-ground core pieces for the remaining 6 dipoles. They report excellent results. They are starting gluing the Purcell Gap sheet pieces down to the core plates. Final assembly of the remaining parts is rapid once the quality of the shim application is proven.

Injector Dipoles (DU/DV)

• The remaining two of three GVs were aligned in the Injector Line.

• The fourth and final injector GV qualified on the test stand (Since it is part of the recirculated line, it will be installed later.

• The remaining two GVs for the extractor region were tested, finding their core fields matched very well for their partnership

Arc 180 Degree Dipoles (GY)

• Wang NMR continued winding the remaining layer of the last GY coil.

• Wang started gluing down the Purcell Gap Shim pieces in anticipation of assembly after potting the coils.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Our goal is to reestablish lasing. We can do that with a completed first arc. With that goal in mind, we decided to temporarily accept the 7 vacuum dipped coils (they are solidly built) and wrap them in Mylar tape to enhance the ground wrap standoff capability. The missing fourth GY coil will be fully vacuum-pressure impregnated. Wang will install them in 4 GQ and 4 GX dipoles and send them to us with no delay. At the time of the third cryomodule installation, we can send them back for full vacuum pressure impregnation that will enhance their longevity.


Trim Quad (QT)

• We qualified four of the 6 newly delivered magnets on the magnet test stand.

Sextupole (SF)

• We drew up sketches of the added pieces needed to transform the original 8 SC Sextupoles into four "Sextupole-Light" magnets. They will have the nomenclature "SG". We received cost estimates for the pieces and modifications from two vendors. At week’s end we were trying to place the order sole source to the company with the much lower estimate rather than wasting time for a bid process that will turn out with the same result.

Octupole (OT)

• The Octupole design was completed by DULY Research.

Beam Line and Vacuum

• We were successful in increasing production rate of the arc chambers at Master Machine. The new procedure is to have Master Machine's welder fit up and tack weld one chamber while another welder does the full edge welding. Using that procedure, the second chamber will be ready Monday and the third is tacked and ready for edge weld. We are indebted to Brian Murphy, a Jefferson Lab Technical Associate, and master welder who offered to be hired to do the edge welding in his off hours. Fortunately, he had the space in his home shop to use a second welding table supplied by Master Machine. The table is instrumental in maintaining the dimensional quality of the chambers.

• X-Chamber - The chamber was installed along with the other pipes in the injection line by the Vacuum Group.

• Y Chamber - It is completed and is off to the vacuum group for leak testing.

• Girder and vacuum assembly continued

WBS 11 (Optics):

We received word from the vendor of the optical cavity vacuum vessels that the two for the IR Upgrade were shipped this past Monday. Receipt is imminent, probably early next week. We will begin mounting UHV-cleaned optical cavity internal parts inside them, so Survey and Alignment can establish fiducials before each assembly is moved into the vault. The Survey and Alignment group has established the mounting coordinates for the outcoupler vacuum vessel and it's associated pedestal and girder for the alignment diagnostics. The corresponding points for the high reflector vacuum vessel and diagnostics will occur next week. The OCMMS hardware has been cleaned for UHV use, with installation next week. We are on schedule to install the

pedestals for the ultraviewers next week.

Hardware to establish a clean environment around the optical cavity vessels was ordered and will be in next week. These will be assembled and installed approximately one week later. I-beams and trolley hoists were ordered - once installed they will be used to lift the vacuum vessel "top hats" on and off the lower section. This will make assembly and testing much easier.

Vacuum outgassing tests established that the vacuum compatible adhesive used by Watlow for attaching their flexible heaters to parts is good into the low 10 E-9 Torr range. The final design review for the optical transport turning mirrors (turning cassettes 1, 3, & 4) was held. Minor corrections will be made to the drawings and they will be checked, then go to Procurement. All parts for the 0F region THz diagnostic should be received by early next week. They will be installed in the X-chamber at the earliest opportunity.

Commissioning activities:

We conducted another test run of the drive laser to test the EPICS controls.


We held a Commissioning Planning Meeting to establish operations ground rules as we bring up the hardware. A number of issues were dealt with; initial decisions were made to assign one CEBAF operator to the team as an interface to future operation under CEBAF Operations control. It was deemed inefficient to have all commissioning carried out by the CEBAF operators since system owners will have to characterize their own systems before they can be handed over for routine operation. Establishment of configuration control was established as a high priority for each system once handed over.

Scheduling of the vault lockup for systems testing will begin in earnest next week. We hope to get a Temporary Operational Safety Procedure for operating the rf systems and beam to 10 MeV signed off next week.